# Direct-to-Cell: A First Look into Starlink’s Direct Satellite-to-Device Radio Access Network through Crowdsourced Measurements

### v6

@@ Abstract
Low Earth Orbit (LEO) satellite mega-constellations have emerged as a viable access solution for broadband connectivity in underserved areas. In 2024, Starlink, in partnership with T-Mobile, began beta testing an SMS-only Supplemental Coverage from Space (SCS) service. This marks the first large-scale deployment of Direct Satellite-to-Device (DS2D) communications, allowing unmodified smartphones to connect directly to spaceborne base stations. This paper presents the first measurement study of deployed DS2D technologies. Using crowdsourced mobile network data from the U.S. between October 2024 and July 2025, we provide evidence-based insights into the capabilities, limitations, and future evolution of DS2D technologies for extending mobile connectivity. We find a strong correlation between the number of satellites deployed, the number of unique cell identifiers measured, and the volume of measurements, concentrated in accessible areas with poor terrestrial network coverage, such as national parks and sparsely populated counties. Stable physical-layer measurements were observed throughout the period, with a $24$-dB lower median RSRP and a $3$-dB higher RSRQ compared to terrestrial networks, reflecting the SMS-only usage of the DS2D network during this period. Based on the SINR measurements collected, we estimate the expected performance of the announced DS2D mobile data service to be around $\mathbf{3}$ Mbps per beam in outdoor conditions. We also discuss strategies to expand this capacity up to $\mathbf{18}$ Mbps in the future, depending on key regulatory and business decisions, including allowable out-of-band emissions, permitted number of satellites, and availability of spectrum and orbital resources.
@@ I Introduction
Following technological developments in fixed broadband connectivity, satellite-based cellular communications, known as Direct Satellite-to-Device (DS2D), have recently emerged as a promising solution to provide seamless connectivity directly to standard unmodified smartphones. Starlink, the satellite communications division of SpaceX, has emerged as the global leader in DS2D communications by rapidly deploying a non-3GPP-NTN-compliant solution that leverages terrestrial International Mobile Telecomunications (IMT) spectrum. In partnership with T-Mobile and using its spectrum, Starlink has been the first company to launch large-scale beta testing of a Supplemental Coverage from Space (SCS) service. The system is initially limited to Short Message Service (SMS), with support for voice and data services planned by late 2025.
In this article, we examine Starlink’s SCS services and characterize its DS2D network—commonly referred to by Starlink as Direct-to-Cell. Using a large-scale dataset provided by the company Weplan Analytics comprising millions of crowdsourced measurements collected from October 2024 to July 2025, our analysis offers valuable insight into the implications of DS2D network design for mobile network planning and spectrum policy. The main contributions of this work are:
  • The first measurement-based study of a commercially deployed DS2D network, including both spatial and temporal analyses.
  • An empirical evaluation of physical-layer Radio Access Network (RAN) parameters, namely Reference Signal Received Power (RSRP), Reference Signal Received Quality (RSRQ) and Signal-to-Interference-plus-Noise Ratio (SINR).
  • An assessment of the projected performance of future data services and strategies for capacity expansion.
@@ II-A Technological Evolution and Emerging Ecosystem
Starlink has pioneered DS2D deployments with a strategy focused on a dense LEO constellation at the lowest feasible orbital altitudes (currently primarily $\sim 550$ km and selectively $\sim 350$ km subject to coordination with the NASA). In contrast, its main competitor, AST SpaceMobile relies on a less dense constellation at higher altitudes ($\sim 700$ km), supported by very large antenna arrays. Although both operators have so far relied primarily on IMT spectrum obtained through agreements with mobile network operators in different countries, they are now pursuing their own terrestrial and satellite spectrum. Other actors in the emerging DS2D ecosystem include Lynk Global, Skylo, and Viasat-Inmarsat. Key use cases comprise ubiquitous connectivity for IoT, supplementary coverage for terrestrial mobile networks, SOS alerts, emergency communications, and maritime and aeronautical connectivity.
@@ II-B Starlink’s Direct Satellite-to-Device (DS2D) Approach
Starlink has reshaped the satellite telecommunication sector by deploying over 7 0007\,000 satellites since its first launches in 2019. The bent-pipe architecture initially adopted [4] enabled rapid deployment using simpler satellite designs at the cost of deploying a dense network of ground stations to ensure continuous coverage. From 2022 onwards, Starlink progressively introduced laser inter-satellite links (ISL), fully adopted in its second-generation (“Gen2”) fleet, enabling onboard signal processing and space-based routing. This transition from a ground-dependent relay to a space-based mesh brings expanded coverage, improved resilience, and enhanced end-to-end performance [1].
Rather than awaiting the completion of 3GPP standardisation of NTN, Starlink pursued an early deployment strategy by integrating conventional LTE eNodeB payloads into its satellites, complemented with proprietary adaptations for NTN operation. This allows smartphones to connect natively using terrestrial mobile standards, while traffic is handled either through the company’s ground station network or, when necessary, routed across the constellation via ISLs [5].
SpaceX announced the completion of its first DS2D constellation in December 2024, consisting of approximately 400 satellites—barely $6\%$ of their deployed fleet. In addition to the SCS service provided in the U.S. through the partnership with T-Mobile, the constellation also offers SCS connectivity in New Zealand and Australia via a partnership with operators and is currently under testing with operators in several other countries.
@@ II-C Related Work
Mirroring the rapid pace of innovation of LEO satellite communication technologies, a growing body of research has emerged from multiple perspectives to explore the design, deployment, and operation of these systems. Early studies focused primarily on the technologies that underpin LEO broadband services, addressing alternative network architectures (e.g., bent-pipe versus ISL designs), hybrid beamforming strategies, interference coordination mechanisms, and resource management challenges [2]. Although much of this work was initially centered on fixed satellite broadband services, recent attention has shifted toward DS2D communications [6], which is reflected in dedicated research and publication initiatives [7]. Alongside theoretical and architectural advances, new simulation tools, such as Hypatia and StarPerf, have been developed to support realistic modeling, design, and performance evaluation of LEO satellite networks under operational conditions.
@@ III Methodology
In this study, we adopt a data-driven, user-centric methodology based on large-scale crowdsourced mobile measurements. The dataset is provided by the crowdsourcing company Weplan Analytics, whose data collection framework is specifically designed to capture and evaluate RAN performance. Crowdsourced measurements offer a unique vantage point for evaluating cellular network performance, capturing real-world user experience even in sparsely populated areas [11]. Subsection III-A describes the data sources and the data preprocessing; Subsection III-B presents how we estimate the potential share of SCS services; and Subsection III-C details how we calculate the expected performance of DS2D communications based on SINR measurements.
@@ III-A Data sources
In addition, we incorporate three complementary data sources that help assess the consistency and reliability of the crowdsourced dataset. First, we compiled a dataset tracking the evolution of the cumulative number of Starlink’s DS2D-capable satellites in orbit by aggregating information from official SpaceX mission reports, which we publish openly as supplementary material. Second, we used the publicly available T-Mobile coverage map to evaluate how well the spatial distribution of DS2D measurements aligns with areas designated as lacking terrestrial coverage. Finally, we used official open-data vector files for U.S. counties and national parks to quantify the SCS share, as detailed next, and official county-level population data from the U.S. Census Bureau to study the correlation of SCS share and population density.
@@ III-B Measuring SCS share
To estimate the likelihood of relying on Starlink SCS services in a particular geographic area, we define a spatial indicator called SCS share. This metric quantifies the relative prevalence of Starlink DS2D connectivity in our dataset by computing the ratio of Starlink DS2D measurements to the total number of measurements (i.e., Starlink DS2D and T-Mobile terrestrial network measurements across all technologies and frequency bands) observed. It thus indicates the degree to which Starlink serves as the primary mobile coverage provider in a given area.
@@ IV Results
This section presents our empirical findings. We start by assessing the relationship between the observed number of cell identifiers and the satellites in orbit in Subsection IV-A. Susbsection IV-B then examines the spatial availability of Starlink’s SCS service, characterizing the evolution of its geographic footprint, and quantifying its SCS share across different types of areas. Subsection IV-C analyzes key RAN performance indicators, including RSRP, RSRQ, and SINR. Lastly, based on these metrics, in Subsection IV-D we estimate the expected capacity of DS2D data services, and discuss and quantify alternatives to improve it.
@@ IV-B Starlink’s DS2D Spatial Analysis
A quantitative analysis using the SCS share indicator (as defined in Section III-B) confirms that large national parks such as Death Valley and Yellowstone exhibit high reliance on DS2D, with values of $0.86$ and $0.80$, respectively. Figure 4 plots the SCS share against population density, computed at a county level. Counties such as Knott, KY; Florence, WI; and Martin, KY, also located near national parks, show SCS share values exceeding $0.85$. Overall, we find a moderately strong negative logarithmic correlation with county-level population density (Pearson correlation coefficient: 0.57-0.57).
@@ IV-C Analysis of DS2D Key Performance Indicators
In this Subsection, we present key metrics of Starlink’s DS2D RAN, namely RSRP, RSRQ and SINR. We examined the dataset for spatial patterns but did not identify any significant trends. In the temporal domain, we observed slight changes. Fig. 5 shows biweekly distributions (in the form of letter-value plots) of RSRP, RSRQ and SINR measurements from the first week of October 2024 (week 41) to the last week of July 2025 (week 32). This timeline captures key milestones of the service setup, including the authorization granted in March 2025 to increase OOBE by $9.4$ dB. The latter, due to transmitter non-linearities, would allow for an approximately $3.1$-dB increase in transmit power according to SpaceX declarations. [^4] See David Goldman’s (Space-X) letter in response to Merissa Velez (FCC) on September 13, 2024. Electronic Comment Filing System (ECFS) code 10914226871971/1. A detailed analysis shows a slight improvement in the distribution of the RAN metrics over the analyzed period, although the observed increment, in particular in RSRP, is smaller than the expected value.
@@
A key question is how this DS2D RAN behavior differs from that based on terrestrial cellular technologies. For this, we benchmark the metrics of this network with those of T-Mobile’s LTE terrestrial network. Fig. 6 compares the cumulative distribution functions (CDF) of RSRP, RSRQ and SINR in Starlink’s DS2D network and T-Mobile terrestrial network. DS2D shows substantially lower and more homogeneous RSRP levels (median: 121–121 dBm vs. 97-97 dBm; interdecile range: $21$ dB vs. $37$ dB), reflecting larger propagation losses and uniform link geometry, as users are located at large but relatively similar distances from the base stations, predominantly with line-of-sight conditions. Despite limited bandwidth, RSRQ is more favorable in DS2D (median: 9-9 dB vs. 12-12 dB) owing to low network load and minimal interference, though this advantage may diminish as demand grows. The correlation between RSRP and RSRQ is stronger in DS2D ($0.69$ vs. $0.34$), consistent with lower interference. SINR values are also lower in DS2D (median: 0 dB vs. $5$ dB), underscoring its coverage-limited, low-load operation.
@@ IV-D Performance assessment of future DS2D services
Since Starlink’s current DS2D operations are restricted to the PCS G Block ($2\times 5$ MHz), the derived spectral efficiency corresponds to an approximate downlink capacity of $3.1$ Mbps per beam. This value should be regarded as an upper-bound estimate of the throughput per connection, corresponding to the scenario of a single user occupying the full bandwidth of the beam under outdoor conditions. While such speeds may be sufficient to provide basic connectivity in remote or underserved regions, they remain well below the average downlink rates offered by terrestrial networks. This gap highlights the complementary nature of DS2D services in their current phase, consistent with their role as an SCS solution.
@@
First, as discussed in Subsection IV-C, we do not observe the expected $3.1$-dB increase in the RAN metrics that would be expected from applyig the FCC’s March 2025 authorization; instead, only a slight improvement is reflected in the measurements, which, with the available evidence, it is not possible to determine whether this variation is attributable to the regulatory change. It therefore remains unclear whether transmissions are already occurring at the newly authorized power levels, although SpaceX’s filings suggest that this is the case (see footnote 4). Nevertheless, in the hypothetical assumption that $9.4$-the dB relaxation in the OOBE limit allows for a future increase in transmit power, and assuming that the increase in SINR would be similar (coverage-limited network), we estimate that the spectral efficiency would increase from the current baseline to $0.89$ bps/Hz (mean) and $0.83$ bps/Hz (median), raising the estimated average per-beam capacity from $3.1$ Mbps to $4.5$ Mbps.
Second, system performance may also be improved by lowering orbital altitudes. SpaceX has been authorized to selectively operate DS2D satellites in Very Low Earth Orbits (VLEO) at an altitude of 340340360360 km, subject to previous coordination with the NASA. [^5] See FCC’s DA 24-1193 Assuming free-space propagation, relocating the entire constellation to lower altitudes would reduce path loss by $20\log_{10}(550/355)\approx 3.8$ dB. To comply with the OOBE constraints, this would require an estimated $1.3$-dB reduction in transmit power to maintain the power flux density levels for OOBE, resulting in a net gain of about $2.5$ dB in received signal power. Assuming again that the SINR values are increased in the same proportion, the spectral efficiency would raise from its baseline value to $0.83$ bps/Hz (mean) and $0.76$ bps/Hz (median). This corresponds to an approximate $1.4\times$ gain, yielding an estimated downlink capacity of $4.2$ Mbps per beam.
@@
Whereas the AWS-4 band has been standardized in 3GPP NTN as n252 and is expected to be integrated into smartphone chipsets in the future, the PCS H Block corresponds to existing IMT spectrum and can therefore be used immediately. Integrating the PCS H Block spectrum into the currently deployed infrastructure, which operates in the PCS G Block, would immediately double system capacity, increasing the per-beam estimate from $3.1$ Mbps to $6.2$ Mbps. In the medium term, once NTN bands are more widely supported in commercial devices, the additional $20$ MHz downlink from AWS-4 could be exploited. Combined with the PCS allocations, this would increase per-beam capacity by a factor of six relative to the baseline, reaching about $18.6$ Mbps when using both Starlink’s spectrum holdings and T-Mobile’s PCS G Block.
Beyond the per-beam capacity expected, these acquisitions reduce Starlink’s reliance on third-party terrestrial operators for spectrum access. By leveraging their own PCS H Block spectrum, Starlink can provide a service comparable to that currently offered through T-Mobile’s spectrum, with the estimated per-beam throughput of 3.1 Mbps. Furthermore, the aggregation of their full spectrum holdings in the AWS and PCS bands (totaling $2\times 25$ MHz) would enable an estimated per-beam throughput of 15.5 Mbps, thereby allowing the provision of DS2D services without reliance on third-party spectrum agreements. Note that these figures correspond to aggregate per-beam capacity rather than achievable single-user rates. Attaining the maximum values per connection would require device support for spectrum aggregation across multiple bands, which is not yet included in current NTN standardization. Accordingly, the realistic upper bound for per-user capacity is expected to be around $12.4$ Mbps (estimated throughput over the AWS band only).
In the longer term, satellite swarm array antennas show promising preliminary results in increasing the directivity of satellite beams [15], which allows for higher gains and improved spatial multiplexing. Achieving this would require the launch and operation of many more satellites, which would also contribute to increasing the overall throughput of DS2D networks and the effective capacity experienced by end users. Although the FCC denied Starlink’s request to authorize 30 00030\,000 DS2D-capable satellites, the company remains authorized to operate up to 7 5007\,500 such satellites, indicating substantial remaining potential for network densification. Furthermore, the aforementioned advancements in network capacity could be combined, potentially enabling greater per-beam throughput in the future.

### v7

@@ Abstract
Low Earth Orbit (LEO) satellite mega-constellations have emerged as a viable access solution for broadband connectivity in underserved areas. In 2024, Starlink, in partnership with T-Mobile, began beta testing an SMS-only Supplemental Coverage from Space (SCS) service. This marks the first large-scale deployment of Direct Satellite-to-Device (DS2D) communications, allowing unmodified smartphones to connect directly to spaceborne base stations. This paper presents the first measurement study of deployed DS2D technologies. Using crowdsourced mobile network data from the U.S. between October 2024 and July 2025, we provide evidence-based insights into the capabilities, limitations, and future evolution of DS2D technologies for extending mobile connectivity. We find a strong correlation between the number of satellites deployed, the number of unique cell identifiers measured, and the volume of measurements, concentrated in accessible areas with poor terrestrial network coverage, such as national parks and sparsely populated counties. Stable physical-layer measurements were observed throughout the period, with a $\mathbf{24}$-dB lower median RSRP and a $\mathbf{3}$-dB higher RSRQ compared to terrestrial networks, reflecting the SMS-only usage of the DS2D network during this period. Based on the SINR measurements collected, we estimate the expected performance of the announced DS2D mobile data service to be around $\mathbf{3}$ Mbps per beam in outdoor conditions. We also discuss strategies to expand this capacity up to $\mathbf{18}$ Mbps in the future, depending on key regulatory and business decisions, including allowable out-of-band emissions, permitted number of satellites, and availability of spectrum and orbital resources.
@@ I Introduction
Following technological developments in fixed broadband connectivity, satellite-based cellular communications, known as Direct Satellite-to-Device (DS2D), have recently emerged as a promising solution to provide seamless connectivity directly to standard unmodified smartphones. Starlink, the satellite communications division of SpaceX, has emerged as the global leader in DS2D communications by rapidly deploying a non-3GPP-NTN-compliant solution that leverages terrestrial International Mobile Telecommunications (IMT) spectrum. In partnership with T-Mobile and using its spectrum, Starlink has been the first company to launch large-scale beta testing of a Supplemental Coverage from Space (SCS) service. The system is initially limited to Short Message Service (SMS), with support for voice and data services planned by late 2025.
In this article, we examine Starlink’s SCS services and characterize its DS2D network—referred to by Starlink as Direct-to-Cell. Using a large-scale dataset provided by the company Weplan Analytics comprising millions of crowdsourced measurements collected from October 2024 to July 2025, our analysis provides valuable insights into the implications of DS2D network design for mobile network planning and spectrum policy. The main contributions of this work are:
  • The first measurement-based study of a commercially operating DS2D network, encompassing both spatial and temporal dimensions.
  • A large-scale empirical evaluation of physical-layer Radio Access Network (RAN) parameters of Starlink’s DS2D network, namely Reference Signal Received Power (RSRP), Reference Signal Received Quality (RSRQ) and Signal-to-Interference-plus-Noise Ratio (SINR).
  • An assessment of the projected performance of Starlink’s future SCS data services and strategies for network capacity expansion.
@@ II-A Technological Evolution and Emerging Ecosystem
Starlink has pioneered DS2D deployments with a strategy focused on a dense LEO constellation at the lowest feasible orbital altitudes (currently primarily $\sim 550$ km and selectively $\sim 350$ km subject to coordination with NASA). In contrast, its main competitor, AST SpaceMobile relies on a less dense constellation at higher altitudes ($\sim 700$ km), supported by very large antenna arrays. Although both operators have so far relied primarily on IMT spectrum obtained through agreements with mobile network operators in different countries, they are now pursuing their own terrestrial and satellite spectrum. Other actors in the emerging DS2D ecosystem include Lynk Global, Skylo, and Viasat-Inmarsat. Key use cases comprise ubiquitous connectivity for IoT, supplementary coverage for terrestrial mobile networks, SOS alerts, emergency communications, and maritime and aeronautical connectivity.
@@ II-B Starlink’s Direct Satellite-to-Device (DS2D) Approach
Starlink has reshaped the satellite telecommunication sector by deploying over 7 0007\,000 satellites since its first launches in 2019. The bent-pipe architecture initially adopted [4] enabled rapid deployment using simpler satellite designs at the cost of deploying a dense network of ground stations to ensure continuous coverage. From 2022 onward, Starlink progressively introduced laser inter-satellite links (ISL), fully adopted in its second-generation (“Gen2”) fleet, enabling onboard signal processing and space-based routing. This transition from a ground-dependent relay to a space-based mesh brings expanded coverage, improved resilience, and enhanced end-to-end performance [1].
Rather than awaiting the completion of 3GPP standardization of NTN, Starlink pursued an early deployment strategy by integrating conventional LTE eNodeB payloads into its satellites, complemented with proprietary adaptations for NTN operation. This allows smartphones to connect natively using terrestrial mobile standards, while traffic is handled either through the company’s ground station network or, when necessary, routed across the constellation via ISLs [5].
SpaceX announced the completion of its first DS2D constellation in December 2024, consisting of approximately 400 satellites—barely $6\%$ of their deployed fleet. In addition to the SCS service provided in the U.S. through the partnership with T-Mobile, the constellation also offers SCS connectivity in New Zealand and Australia via a partnership with operators and is under testing with operators in several other countries.
@@ II-C Related Work
Mirroring the rapid pace of innovation of LEO satellite communication technologies, a growing body of research has emerged from multiple perspectives to explore the design, deployment, and operation of these systems. Early studies focused primarily on the technologies that underpin LEO broadband services, addressing alternative network architectures (e.g., bent-pipe versus ISL designs), hybrid beamforming strategies, interference coordination mechanisms, and resource management challenges [2]. Although much of this work was initially centered on fixed satellite broadband services, recent attention has shifted toward DS2D communications [6], which is reflected in dedicated research and publication initiatives. [^4] See the Special Issue Direct Satellite-to-Device Communications: Technologies, Connectivity, and Spectrum Management of the IEEE Communications Magazine (2025). Alongside theoretical and architectural advances, new simulation tools, such as Hypatia and StarPerf, have been developed to support realistic modeling, design, and performance evaluation of LEO satellite networks under operational conditions.
@@ III Methodology
In this study, we adopt a data-driven, user-centric methodology based on large-scale crowdsourced mobile measurements. Crowdsourced measurements offer a unique vantage point for evaluating cellular network performance, capturing real-world user experience even in sparsely populated areas [10]. The dataset is provided by the crowdsourcing company Weplan Analytics, whose data collection framework is specifically designed to capture and evaluate RAN performance. While the raw crowdsourced measurements are proprietary and cannot be shared, the derived statistics and analytical procedures are fully described in this paper to enable methodological reproducibility. Subsection III-A describes the data sources and the data preprocessing; Subsection III-B presents how we estimate the potential share of SCS services; and Subsection III-C details how we calculate the expected performance of DS2D communications based on SINR measurements.
@@ III-A Data sources
In addition, we incorporate three complementary data sources that help assess the consistency and reliability of the crowdsourced dataset. First, we compiled a dataset tracking the evolution of the cumulative number of Starlink’s DS2D-capable satellites in orbit by aggregating information from official SpaceX mission reports, which we publish openly as supplementary material [12]. Second, we used the publicly available T-Mobile coverage map to evaluate how well the spatial distribution of DS2D measurements aligns with areas designated as lacking terrestrial coverage. Finally, we used official open-data vector files for U.S. counties and national parks to quantify the SCS share, as detailed next, and official county-level population data from the U.S. Census Bureau to study the correlation of SCS share and population density.
@@ III-B Measuring SCS share
To estimate the likelihood of relying on Starlink SCS services in a particular geographic area, we define a spatial indicator called SCS share. This metric quantifies the relative prevalence of Starlink DS2D connectivity in our dataset by computing the ratio of Starlink DS2D measurements to the total number of measurements observed (i.e., Starlink DS2D and T-Mobile terrestrial network measurements across all technologies and frequency bands). It thus indicates the degree to which Starlink serves as the primary mobile coverage provider in a given area.
@@ IV Results
This section presents our empirical findings. We start by assessing the relationship between the observed number of cell identifiers and the satellites in orbit in Subsection IV-A. Subsection IV-B then examines the spatial availability of Starlink’s SCS service, characterizing the evolution of its geographic footprint, and quantifying its SCS share across different types of areas. Subsection IV-C analyzes key RAN performance indicators, including RSRP, RSRQ, and SINR. Lastly, based on these metrics, in Subsection IV-D we estimate the expected capacity of DS2D data services, and discuss and quantify alternatives to improve it.
@@ IV-B Starlink’s DS2D Spatial Analysis
A quantitative analysis using the SCS share indicator (as defined in Section III-B) confirms that large national parks such as Death Valley and Yellowstone exhibit high reliance on DS2D, with values of $0.86$ and $0.80$, respectively. Figure 4 plots the SCS share against population density, computed at a county level. Counties such as Knott (KY). Florence (WI), and Martin (KY), also located near national parks, show SCS share values exceeding $0.85$. Overall, we find a moderately strong negative logarithmic correlation with county-level population density (Pearson correlation coefficient: 0.57-0.57).
@@ IV-C Analysis of DS2D Key Performance Indicators
In this Subsection, we present key metrics of Starlink’s DS2D RAN, namely RSRP, RSRQ and SINR. We examined the dataset for spatial patterns but did not identify any significant trends. In the temporal domain, we observed slight changes. Fig. 5 shows biweekly distributions (in the form of letter-value plots) of RSRP, RSRQ and SINR measurements from the first week of October 2024 (week 41) to the last week of July 2025 (week 32). This timeline captures key milestones of the service setup, including the authorization granted in March 2025 to increase OOBE by $9.4$ dB. The latter, due to transmitter non-linearities, would allow for an approximately $3.1$-dB increase in transmit power according to SpaceX declarations. [^5] See David Goldman’s (Space-X) letter in response to Merissa Velez (FCC) on September 13, 2024. Electronic Comment Filing System (ECFS) code 10914226871971/1 A detailed analysis shows a slight improvement in the distribution of the RAN metrics over the analyzed period, although the observed increment, in particular in RSRP, is smaller than the expected value.
@@
A key question is how this DS2D RAN behavior differs from that of terrestrial cellular technologies. For this, we benchmark the metrics of this network with those of T-Mobile’s LTE terrestrial network. Fig. 6 compares the cumulative distribution functions (CDF) of RSRP, RSRQ and SINR in Starlink’s DS2D network and T-Mobile terrestrial network. DS2D shows substantially lower and more homogeneous RSRP levels (median: 121–121 dBm vs. 97-97 dBm; interdecile range: $21$ dB vs. $37$ dB), reflecting larger propagation losses and uniform link geometry, as users are located at large but relatively similar distances from the base stations, predominantly with line-of-sight conditions. Despite limited bandwidth, RSRQ is more favorable in DS2D (median: 9-9 dB vs. 12-12 dB) owing to low network load and minimal interference, though this advantage may diminish as demand grows. The correlation between RSRP and RSRQ is stronger in DS2D ($0.69$ vs. $0.34$), consistent with lower interference. SINR values are also lower in DS2D (median: 0 dB vs. $5$ dB), underscoring its coverage-limited, low-load operation.
@@ IV-D Performance assessment of future DS2D services
Since Starlink’s current DS2D operations are restricted to the PCS G Block ($2\times 5$ MHz), the derived spectral efficiency corresponds to an approximate downlink capacity of $3.1$ Mbps per beam. This value should be regarded as an upper-bound estimate of the throughput per connection, corresponding to the scenario of a single user occupying the full bandwidth of the beam. While such speeds may be sufficient to provide basic connectivity in remote or underserved regions, they remain well below the average downlink rates offered by terrestrial networks. This gap highlights the complementary nature of DS2D services in their current phase, consistent with their role as an SCS solution.
@@
First, as discussed in Subsection IV-C, we do not observe the expected $3.1$-dB increase in the RAN metrics that would be expected from applying the FCC’s March 2025 authorization; instead, only a slight improvement is reflected in the measurements, which, with the available evidence, it is not possible to determine whether this variation is attributable to the regulatory change. It therefore remains unclear whether transmissions are already occurring at the newly authorized power levels, although SpaceX’s filings suggest that this is the case (see footnote 4). Nevertheless, in the hypothetical assumption that the $9.4$-dB relaxation in the OOBE limit allows for a future increase in transmit power, and assuming that the increase in SINR would be similar (coverage-limited network), we estimate that the spectral efficiency would increase from the current baseline to $0.89$ bps/Hz (mean) and $0.83$ bps/Hz (median), raising the estimated average per-beam capacity from $3.1$ Mbps to $4.5$ Mbps.
Second, system performance may also be improved by lowering orbital altitudes. SpaceX has been authorized to selectively operate DS2D satellites in Very Low Earth Orbits (VLEO) at an altitude of 340340360360 km, subject to previous coordination with NASA. [^6] See FCC’s DA 24-1193 Assuming free-space propagation, relocating the entire constellation to lower altitudes would reduce path loss by $20\log_{10}(550/355)\approx 3.8$ dB. To comply with the OOBE constraints, this would require an estimated $1.3$-dB reduction in transmit power to maintain the power flux density levels for OOBE, resulting in a net gain of about $2.5$ dB in received signal power. Assuming again that the SINR values are increased in the same proportion, the spectral efficiency would raise from its baseline value to $0.83$ bps/Hz (mean) and $0.76$ bps/Hz (median). This corresponds to an approximate $1.4\times$ gain, yielding an estimated downlink capacity of $4.2$ Mbps per beam.
@@
Whereas the AWS-4 band has been standardized in 3GPP NTN as n252 and is expected to be integrated into smartphone chipsets in the future, the PCS H Block corresponds to existing IMT spectrum and can therefore be used immediately. Integrating the PCS H Block spectrum into the currently deployed infrastructure—which operates in the PCS G Block—would immediately double system capacity, increasing the per-beam estimate from $3.1$ Mbps to $6.2$ Mbps. In the medium term, once NTN bands are more widely supported in commercial devices, the additional $20$ MHz downlink from AWS-4 could be exploited. Combined with the PCS allocations, this would increase per-beam capacity by a factor of six relative to the baseline, reaching about $18.6$ Mbps when using both Starlink’s spectrum holdings and T-Mobile’s PCS G Block.
Beyond the per-beam capacity expected, these acquisitions reduce Starlink’s reliance on third-party terrestrial operators for spectrum access. By leveraging their own PCS H Block spectrum, Starlink can provide a service comparable to that currently offered through T-Mobile’s spectrum, with the estimated per-beam throughput of 3.1 Mbps. Furthermore, the aggregation of their full spectrum holdings in the AWS and PCS bands (totaling $2\times 25$ MHz) would enable an estimated per-beam throughput of 15.5 Mbps, thereby allowing the provision of DS2D services without reliance on third-party spectrum agreements. Note that these figures correspond to aggregate per-beam capacity rather than achievable single-user rates. Attaining the maximum values per connection would require device support for spectrum aggregation across these multiple bands, which is not yet included in current NTN standardization. Accordingly, the realistic upper bound for per-user capacity is expected to be around $12.4$ Mbps (estimated throughput over the AWS band only).
In the longer term, satellite swarm antenna arrays show promising preliminary results in increasing the directivity of satellite beams [15], which allows for higher gains and improved spatial multiplexing. Achieving this would require the launch and operation of many more satellites, which might also contribute to increasing the overall throughput of DS2D networks and the effective capacity experienced by end users. Although the FCC denied Starlink’s request to authorize 30 00030\,000 DS2D-capable satellites, the company remains authorized to operate up to 7 5007\,500 such satellites, indicating substantial remaining potential for network densification. Furthermore, the aforementioned advancements in network capacity could be combined, potentially enabling greater per-beam throughput in the future.